# Cluster state as a non-invertible symmetry protected topological phase

### v1

@@ II Non-invertible symmetry of the cluster model
The operators $\eta^{\mathrm{e}}$, $\eta^{\mathrm{o}}$, and $\mathsf{D}$ commutes with the Hamiltonian and satisfy the following algebra
@@
We have an (internal) non-invertible symmetry in the cluster model that does not mix with the lattice translation. Therefore, the cluster state is a topological phase protected by a non-invertible symmetry.
In Appendix B.3, we show that this non-invertible symmetry is described by the fusion category Rep(D8), whose fusion algebra is given by the tensor product of the irreducible representations of the group D8. (See [34] for a review of fusion category, which is built on [35, 36].) We prove this by gauging the non-invertible symmetry and find a dual D8 symmetry.
@@ III.2 Three Rep(D8) SPT phases
|$\ket{\text{odd}}=\prod_{n=1}^{L/2}Z_{2n}Z_{2n+2}\,e^{\frac{\pi i}{4}(Z_{2n-1}Z_{2n+1}-1)}\ket{\text{cluster}}\,.$ (20)
@@ IV Edge modes
In the previous section, we distinguished the three SPT phases by their different symmetry breaking patterns after gauging a part of the non-invertible symmetry. Here, we discuss the edge modes on the boundary of these phases. For ordinary on-site symmetries, non-trivial edge modes appear on the boundary between a (nontrivial) SPT phase and the product state. As emphasized before, there is no canonical notion of a trivial SPT phase for non-invertible symmetries, and we are instructed to identify the edge modes at the interface between two distinct SPT phases.
@@
In general, the three SPT phases are distinguished by the projective algebra involving the local factors of $\mathsf{D}$ and $\eta^{\mathrm{e}},\eta^{\mathrm{o}}$. The projective algebra implies that the defect of the non-invertible symmetry $\mathsf{D}$ carries a non-trivial charge under $\eta^{\mathrm{e}},\eta^{\mathrm{o}}$, or $\eta^{\mathrm{d}}$, which will be explored in [43]. In the continuum, these charges are denoted by $\nu(\mathrm{e}),\nu(\mathrm{o}),\nu(\mathrm{d})$ in [33, 26, 38]. Their values for the three SPT phases are [^10] These invariants explain our terminology of ‘odd’ and ‘even’ SPT phases and suggest that the cluster state can be called the ‘diagonal’ SPT.
@@ V No symmetric entangler for non-invertible SPT phases
We now argue for the more general statement without assuming $\mathsf{U}$ is Clifford. The idea it to show that if such a (unitary) symmetric entangler $\mathsf{U}$ exists, then there must exist a unitary operator $\hat{\mathsf{U}}$ in the $\mathbf{TST}$ gauged/KT transformed system such that it is invariant under the dual symmetry ^$\hat{{\mathbb{Z}}}_{2}^{\mathrm{e}}\times\hat{{\mathbb{Z}}}_{2}^{\mathrm{o}}\times\hat{{\mathbb{Z}}}_{2}^{\mathsf{V}}$ and maps the two symmetry breaking phases to each other. However, this is impossible since the two SPT phases correspond to two different patterns of symmetry breaking after gauging and hence cannot be related by a symmetric unitary operator.
@@ B.2 Type III anomaly and the Levin-Gu model
where $\hat{A}^{\mathrm{e}},\hat{A}^{\mathrm{o}},\hat{A}^{\mathsf{V}}$ are background ${\mathbb{Z}}_{2}$ gauge fields. This is referred to as a type III anomaly [47, 48] because it involves all three generators of ^$\hat{\mathbb{Z}}_{2}^{\mathrm{e}}\times\hat{\mathbb{Z}}_{2}^{\mathrm{o}}\times\hat{\mathbb{Z}}_{2}^{\mathsf{V}}$. As a consistency check, this reduces to the Levin-Gu anomaly for the diagonal symmetry by setting $\hat{A}^{\mathrm{e}}=\hat{A}^{\mathrm{o}}=\hat{A}^{\mathsf{V}}$. One feature of this type III anomaly is that out of the 7 ${\mathbb{Z}}_{2}$ subgroup of ^$\hat{\mathbb{Z}}_{2}^{\mathrm{e}}\times\hat{\mathbb{Z}}_{2}^{\mathrm{o}}\times\hat{\mathbb{Z}}_{2}^{\mathsf{V}}$, only the one generated by $\hat{\eta}^{\mathrm{e}}\hat{\eta}^{\mathrm{o}}\hat{\mathsf{V}}$ has an anomaly.
@@
To summarize, the twisted gauging maps a non-invertible symmetry Rep(D8) to an anomalous invertible symmetry ^$\hat{\mathbb{Z}}_{2}^{\mathrm{e}}\times\hat{\mathbb{Z}}_{2}^{\mathrm{o}}\times\hat{\mathbb{Z}}_{2}^{V}$:
[…]
@@ B.3 Gauging Rep($\mathrm{D}_{8}$) in steps
$-\sum_{j=1}^{L}X_{j}\left(\frac{1+\tilde{Z}_{j}}{2}(1+\hat{Z}_{j-1}\hat{Z}_{j+1})+\frac{1-\tilde{Z}_{j}}{2}(\hat{Z}_{j-1}+\hat{Z}_{j+1})\right)$
@@ E.2 $\mathsf{D}\ket{\text{odd}}=2(-1)^{L/4}\ket{\text{odd}}$
$U_{\text{odd}}=\prod_{n=1}^{L/2}Z_{2n}Z_{2n+2}\,e^{\frac{\pi i}{4}(Z_{2n-1}Z_{2n+1}-1)}\,.$ (81)

### v2

@@ II Non-invertible symmetry of the cluster model
The operators $\eta^{\mathrm{e}}$, $\eta^{\mathrm{o}}$, and $\mathsf{D}$ commute with the Hamiltonian and satisfy the following algebra
@@
Note that even though the definition of $\mathsf{D}$ appears to involve a lattice translation $T$, the factors $\mathsf{D}^{\mathrm{e}},\mathsf{D}^{\mathrm{o}}$ each involves a half translation in the opposite direction, and hence the algebra of $\mathsf{D}$ does not mix with lattice translations. We conclude that the cluster state is a topological phase protected by a non-invertible symmetry.
In Appendix B.3, we show that this non-invertible symmetry is described by the fusion category Rep(D8), whose fusion algebra is given by the tensor product of the irreducible representations of the group D8. (See [34] for a review of fusion category, which is built on [35, 36].) We prove this by gauging the non-invertible symmetry and finding a dual D8 symmetry.
@@ III.2 Three Rep(D8) SPT phases
|$\ket{\text{odd}}=\prod_{k=1}^{L/4}Z_{4k-2}Z_{4k}\prod_{n=1}^{L/2}e^{\frac{\pi i}{4}(Z_{2n-1}Z_{2n+1}-1)}\ket{\text{cluster}}\,.$ (20)
@@ IV Edge modes
In the previous section, we distinguished the three SPT phases by their different symmetry breaking patterns after gauging a part of the non-invertible symmetry. Here, we discuss the edge modes on the boundary of these phases.
For ordinary on-site symmetries, a (non-trivial) SPT phase is characterized by the edge modes or the projective algebra of the symmetry operator on an open chain with two boundaries. However, our non-invertible symmetry is not a product of local unitary operators, and it is unclear how to “cut it open.” We therefore need a more universal diagnostic for SPT phases, one that is equivalent to the standard one for on-site symmetries, but also applicable to symmetries that are not necessarily on-site. To this end, we reinterpret an open chain system as an interface between the (non-trivial) SPT phase and the product state on a closed chain. The relative difference between this SPT phase and the product state is then diagnosed by the edge modes at the interfaces. The advantage of this simple reinterpretation is that we never have to cut open the symmetry operator, and this criterion can be readily generalized to non-invertible symmetries. See Appendix D for a detailed discussion of this reinterpretation of the edge modes in the cluster model.
@@
In general, the three SPT phases are distinguished by the projective algebra involving the local factors of $\mathsf{D}$ and $\eta^{\mathrm{e}},\eta^{\mathrm{o}}$. The projective algebra implies that the defect of the non-invertible symmetry $\mathsf{D}$ carries a non-trivial charge under $\eta^{\mathrm{e}},\eta^{\mathrm{o}}$, or $\eta^{\mathrm{d}}$, which will be explored elsewhere. In the continuum, these charges are denoted by $\nu(\mathrm{e}),\nu(\mathrm{o}),\nu(\mathrm{d})$ in [33, 26, 38]. Their values for the three SPT phases are [^10] These invariants explain our terminology of ‘odd’ and ‘even’ SPT phases and suggest that the cluster state can be called the ‘diagonal’ SPT.
@@ V No symmetric entangler for non-invertible SPT phases
We now argue for the more general statement without assuming $\mathsf{U}$ is Clifford. The idea is to show that if such a (unitary) symmetric entangler $\mathsf{U}$ exists, then there must exist a unitary operator $\hat{\mathsf{U}}$ in the $\mathbf{TST}$ gauged/KT transformed system such that it is invariant under the dual symmetry ^$\hat{{\mathbb{Z}}}_{2}^{\mathrm{e}}\times\hat{{\mathbb{Z}}}_{2}^{\mathrm{o}}\times\hat{{\mathbb{Z}}}_{2}^{\mathsf{V}}$ and maps the two symmetry breaking phases to each other. However, this is impossible since the two SPT phases correspond to two different patterns of symmetry breaking after gauging and hence cannot be related by a symmetric unitary operator.
@@ B.2 Type III anomaly and the Levin-Gu model
where $\hat{A}^{\mathrm{e}},\hat{A}^{\mathrm{o}},\hat{A}^{\mathsf{V}}$ are background ${\mathbb{Z}}_{2}$ gauge fields. This is referred to as a type III anomaly [47, 48] because it involves all three generators of ^$\hat{\mathbb{Z}}_{2}^{\mathrm{e}}\times\hat{\mathbb{Z}}_{2}^{\mathrm{o}}\times\hat{\mathbb{Z}}_{2}^{\mathsf{V}}$. As a consistency check, this reduces to the Levin-Gu anomaly for the diagonal symmetry by setting $\hat{A}^{\mathrm{e}}=\hat{A}^{\mathrm{o}}=\hat{A}^{\mathsf{V}}$. One feature of this type III anomaly is that out of the 7 ${\mathbb{Z}}_{2}$ subgroups of ^$\hat{\mathbb{Z}}_{2}^{\mathrm{e}}\times\hat{\mathbb{Z}}_{2}^{\mathrm{o}}\times\hat{\mathbb{Z}}_{2}^{\mathsf{V}}$, only the one generated by $\hat{\eta}^{\mathrm{e}}\hat{\eta}^{\mathrm{o}}\hat{\mathsf{V}}$ has an anomaly.
@@ B.3 Gauging Rep($\mathrm{D}_{8}$) in steps
$-\sum_{j=1}^{L}\hat{X}_{j}\left(\frac{1+\tilde{Z}_{j}}{2}(1+\hat{Z}_{j-1}\hat{Z}_{j+1})+\frac{1-\tilde{Z}_{j}}{2}(\hat{Z}_{j-1}+\hat{Z}_{j+1})\right)$
@@ E.2 $\mathsf{D}\ket{\text{odd}}=2(-1)^{L/4}\ket{\text{odd}}$
$U_{\text{odd}}=\prod_{n=1}^{L/2}Z_{2n}\,e^{\frac{\pi i}{4}(Z_{2n-1}Z_{2n+1}-1)}\,.$ (80)